# StarFlow: Leveraging Normalizing Flows for Stellar Age Estimation in SDSS-V DR19

### v1

@@ 1 Introduction
Understanding the ages of stars is crucial for unraveling the formation history and evolution of our Galaxy. The Milky Way offers a unique opportunity to study galactic formation up close by resolving individual stars and directly measuring their properties. While extragalactic studies beyond the local group rely on integrated light and population synthesis models, the detailed observations within our own Galaxy allow for a more precise reconstruction of its chemodynamical evolution. Spectroscopic surveys such as SDSS-IV/APOGEE (Majewski et al., 2017; Wilson et al., 2019, Apache Point Observatory Galactic Evolution Experiment;), SDSS-V/MWM (Smee et al., 2013; Kollmeier et al., 2019, Milky Way ;, Kollmeier et al. (2025) in preparation), GALactic Archaeology with HERMES (GALAH; Buder et al., 2018), Large Sky Area Multi-Object Fiber Spectroscopic Telescope (LAMOST; Cui et al., 2012), and Gaia-ESO (Gilmore et al., 2012), along with astrometric data from Gaia (Gaia Collaboration et al., 2016), have provided unprecedented insights into the chemical abundances, kinematics, and dynamics of millions of stars. However, accurately determining stellar ages remains one of the most challenging tasks in astrophysics because age is not a directly observable property (e.g., Soderblom, 2010).
Determining stellar ages is especially challenging because it must be inferred from other measurable quantities, leading to degeneracies and uncertainties. Traditional methods like isochrone fitting compare observed stellar parameters—such as effective temperature and luminosity—to theoretical models to estimate ages (Serenelli et al., 2013; Pont & Eyer, 2004). This technique is effective for certain evolutionary phases, such as stars on the sub-giant branch, where evolutionary tracks in the Hertzsprung-Russell diagram are well-separated. However, for more evolved stars, evolutionary tracks converge, resulting in degenerate solutions and increased uncertainties in age estimation. Observational uncertainties further compound these difficulties, making precise age determination problematic.
@@
An alternative method involves using the carbon-to-nitrogen abundance ratio ([C/N]), which changes during the red giant branch phase due to internal mixing processes. As stars ascend the giant branch, their convective layers dredge up material processed in the carbon-nitrogen-oxygen (CNO) cycle, altering the surface [C/N] ratio in a way that correlates with stellar mass—and thus age (Martig et al., 2016b). However, modeling these mixing processes accurately is complex, and uncertainties in stellar models make it difficult to predict [C/N] ratios precisely (Masseron & Gilmore, 2015; Roberts et al., 2024). Furthermore, stars with $[\rm{Fe/H}]<-0.5$ experience extra-mixing processes, complicating the [C/N] - age relationship (Shetrone et al., 2019). These limitations makes empirical calibration, such as using asteroseismic masses to calibrate the [C/N]-mass relation, an attractive option.
@@ 4.1 Model Training and Optimization
The augmented dataset then undergoes standardization, where each parameter’s mean is subtracted and then divided by its standard deviation. This standardized data is subsequently split into training, validation, and test sets with ratios of 0.7, 0.25, and 0.05, respectively. A blank model, as described in Section 2.3, is initialized with PyTorch’s default uniform initialization (i.e., weights drawn from $[-1/\sqrt{n},1/\sqrt{n}]$), and the augmented and standardized training dataset is organized into batches of $2^{16}$ (65,536) samples, which are used directly in each training epoch. During each epoch, the model computes the forward KL loss for the current training batch as well as for the validation set. The Adam optimizer is then used to update the model parameters based on the training batch loss, and the loss from the validation set is recorded. This validation loss serves as a metric to monitor training progress using data that are not included in the training set. Training is halted once the validation loss ceases to decrease for 100 consecutive epochs. Our model took $\approx 80,000$ epochs before ceasing.
@@ 4.2 Deriving Age Posteriors from Normalizing Flows
Figure 2 shows an example age posterior recovery for a star in the test set. The relation between the orange curve and the blue histogram shows how the recovered age posterior approximates the underlying distribution of ages in the training set. The histogram does not account for input uncertainties in the data, however the recovered posterior shown in orange does. This example happens to just barely succeed in recovering the age to $1\sigma$ and demonstrates how the posterior from normalizing flows can have asymmetric uncertainties.
@@
Figure 2: (Left) The larger black point marks the location of an example test-set star overlaid on the full training set (color-coded by [Fe/H]) shown in a Kiel diagram (Upper Left) and in the [C/N]–age relation (Lower Left). In the [C/N]–age plot, a red error bar indicates the model’s age estimate and uncertainty for the star. (Right) Age posterior for the example star. The black curve represents the full posterior computed with the input uncertainties, with the solid orange line indicating the maximum likelihood and dashed lines showing $1\sigma$ error bars. The red dashed line indicates the asteroseismic age, while the blue histogram displays the age distribution of training-set stars that are nearby in input parameter space.
@@ 4.4 Model Predictions Beyond Stellar Ages
Figure 3: (Left) 2D probability distribution of [C/Fe] & [N/Fe] for the star depicted in Fig 2. The red dot marks the maximum likelihood, and the red line indicates the $1\sigma$ boundary. The orange point and error bar shows the star’s measured [C/Fe] & [N/Fe] and associated uncertainties. The diagonals show lines of constant [C/N]. (Right) Displays a similar plot for a different star with a higher [C/Fe] than the model predicts. Both plot exhibit slight banding features, these are most likely artifacts of the normalizing flow model over fitting to local noise or small-scale features.
@@ 4.5 Training Space Density
Figure 4: Parameter space coverage of stars with a training space density above $3\times 10^{9}$ (color-coded by density) compared to the full DR19 dataset (black points). These are scatter plots with each stars training density used for color coding. Regions with higher densities are better sampled in the training data, yielding more reliable age estimates. (Left) Kiel diagram (Right) [Fe/H] versus surface gravity ($\log g$). The sharp banded features seen in the black points are artifacts from the ASPCAP grid edges. This Figure is repeated in Appendix B.1 with different training space density thresholds.
We utilize this value as a measure of how well-sampled the parameter space around a given star’s parameters was during training, and we use it as a cutoff to limit age estimates to stars that occupy well-covered regions of parameter space. Figure 4 shows a Kiel diagram of the stars in our catalog, color-coded by their training space density. The right-hand panel shows the same sample but restricted to stars with a training space density greater than $3\times 10^{9}$, closely reflecting the parameter space coverage of the original training set. Choosing a smaller cutoff allows the model to extrapolate into less well-sampled regions, while a higher cutoff restricts the sample to stars within the most densely populated parts of parameter space. In subsequent sections, we refer to this value as the training space density, and we adopt a cutoff of greater than $3\times 10^{9}$ for the age validation. This choice demonstrates confidence in the model’s ability to extrapolate slightly beyond the asteroseismic core of the training data. We feel that this adopted value is a good starting point that is not too conservative, resulting in a small number of stars in the output, while also constraining the model from extrapolating too far. The implications of this choice on the use of our ages for MWM are discussed in Section 6.1. Appendix B shows how varying the training space density used as a threshold changes the various plots and maps shown in this work.
@@ 5.1 Validation Against Asteroseismic ages
Figure 5 shows the maximum likelihood ages for the test set as determined by our model compared to the asteroseismic ages. The figure demonstrates that for ages $<8$ Gyr the model returns accurate ages. However, at ages $>8$ Gyr, the model’s maximum likelihood begins to diverge from the asteroseismic estimates. This result is not unexpected, as it has been observed and documented in multiple spectroscopic age studies (Leung & Bovy 2019; Mackereth et al. 2019; Anders et al. 2023; Stone-Martinez et al. 2024). The plateau at 8 Gyr is also expected from [C/N] ages, as shown in Roberts et al. (2024), due to mixing processes being weaker in lower mass (higher age) stars.
@@ 5.2 Evaluating Model-Derived Uncertainties
Figure 7: (Left) Histogram of the relative $1\sigma$ uncertainties for asteroseismic labels (blue), the averaged upper and lower relative $1\sigma$ uncertainties derived from the StarFlow model (orange), and the relative differences between the asteroseismic age and the StarFlow maximum likelihood ages (green). (Right) Bar plot showing the median relative uncertainty recovered from the StarFlow age posteriors, binned by asteroseismic age.
@@ Acknowledgments
We are grateful to Juna Kollmeier, Jamie Tayar, and José G. Fernández-Trincado for valuable discussions and feedback. Much of the discussion in this manuscript was inspired by conversations with various conference and seminar attendees, including Phil Van-Lane, James Johnson, Jennifer Johnson, Marc Pinsonneault, Jason Jackiewicz, David Weinberg, Phillip Cargile, Henry Leung, Joshua Speagle, and Paul Beck. A.S-M. gratefully acknowledges support from the SDSS-V Data Inspection program. E.J.G. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2202135 Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org. SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including the Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, Caltech, the Gotham Participation Group, Harvard University, Heidelberg University, The Flatiron Institute, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University. This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix B Using a different training density threshold
Here we present how Figures 4, 10, and 13 change when varying the training density threshold used.
@@
B.1 Figure 4: Parameter space coverage
@@
B.2 Figure 10: Galactic Age Map
@@
B.3 Figure 13: Chemo-Age map

### v2

@@ 1 Introduction
Understanding the ages of stars is crucial for unraveling the formation history and evolution of our Galaxy. The Milky Way offers a unique opportunity to study galactic formation up close by resolving individual stars and directly measuring their properties. While extragalactic studies beyond the local group rely on integrated light and population synthesis models, the detailed observations within our own Galaxy allow for a more precise reconstruction of its chemodynamical evolution. Spectroscopic surveys such as SDSS-IV/APOGEE (Majewski et al., 2017; Wilson et al., 2019, Apache Point Observatory Galactic Evolution Experiment;), SDSS-V/MWM (Smee et al., 2013; Kollmeier et al., 2019, Milky Way Mapper;, Kollmeier et al. (2025) in preparation), GALAH (Buder et al., 2018; De Silva et al., 2015, GALactic Archaeology with HERMES;), LAMOST (Cui et al., 2012; Deng et al., 2012, Large Sky Area Multi-Object Fiber Spectroscopic Telescope;), and Gaia-ESO (Gilmore et al., 2012), along with astrometric data from Gaia (Gaia Collaboration et al., 2016), have provided unprecedented insights into the chemical abundances, kinematics, and dynamics of millions of stars. However, accurately determining stellar ages remains one of the most challenging tasks in astrophysics because age is not a directly observable property (e.g., Soderblom, 2010).
Determining stellar ages is especially challenging because it must be inferred from other measurable quantities, leading to degeneracies and uncertainties. Traditional methods like isochrone fitting compare observed stellar parameters—such as effective temperature and luminosity—to theoretical models to estimate ages (Pont & Eyer, 2004; Dotter et al., 2008; Bressan et al., 2012; Serenelli et al., 2013). This technique is effective for certain evolutionary phases, such as stars on the sub-giant branch, where evolutionary tracks in the Hertzsprung-Russell diagram are well-separated. However, for more evolved stars, evolutionary tracks converge, resulting in degenerate solutions and increased uncertainties in age estimation. Observational uncertainties further compound these difficulties, making precise age determination problematic.
@@
An alternative method involves using chemical clocks. The premise is that certain elemental abundances trace stellar age, usually [Fe/H], [C/N], [$\alpha$/Fe], or other elements such as R-process elements. [$\alpha$/Fe], [Fe/H] and R-process abundance based ages are predicted in classical chemical evolution models Matteucci & Recchi, 2001; Pagel, 2009; Nissen, 2015, e.g., and observed in the solar neighborhood (Fuhrmann, 2011; Haywood, Misha et al., 2013). However, using ages estimated from these elements can introduce biases when examining chemical evolution due their calibration being dependent on current understanding of Galactic chemical evolution. In contrast, [C/N] is tied directly to stellar evolution (Masseron & Gilmore, 2015; Martig et al., 2016b; Roberts et al., 2024). The carbon-to-nitrogen abundance ratio ([C/N]) changes during the red giant branch phase due to internal mixing processes. As stars ascend the giant branch, their convective layers dredge up material processed in the carbon-nitrogen-oxygen (CNO) cycle, altering the surface [C/N] ratio in a way that correlates with stellar mass—and thus age (Martig et al., 2016b). However, modeling these mixing processes accurately is complex, and uncertainties in stellar models make it difficult to predict [C/N] ratios precisely (Masseron & Gilmore, 2015; Roberts et al., 2024). Furthermore, stars with $[\rm{Fe/H}]<-0.5$ experience extra-mixing processes, complicating the [C/N]-age relationship (Shetrone et al., 2019). These limitations make empirical calibration, such as using asteroseismic masses to calibrate the [C/N]-mass relation, an attractive option.
@@ 4.1 Model Training and Optimization
The augmented dataset then undergoes standardization, where each parameter’s mean is subtracted and then divided by its standard deviation. This standardized data is subsequently split into training, validation, and test sets with ratios of 0.7, 0.25, and 0.05, respectively. A blank model, as described in Section 2.3, is initialized with PyTorch’s default uniform initialization (i.e., weights drawn from $[-1/\sqrt{n},1/\sqrt{n}]$), and the augmented and standardized training dataset is organized into batches of $2^{16}$ (65,536) samples, which are used directly in each training epoch. No weight adjustments are applied to any of the training stars. During each epoch, the model computes the forward KL loss for the current training batch as well as for the validation set. The Adam optimizer is then used to update the model parameters based on the training batch loss, and the loss from the validation set is recorded. This validation loss serves as a metric to monitor training progress using data that are not included in the training set. Training is halted once the validation loss ceases to decrease for 100 consecutive epochs. Our model took $\approx 80,000$ epochs before ceasing.
@@ 4.2 Deriving Age Posteriors from Normalizing Flows
Figure 2 shows a representative test-set star with asteroseismic age near the older end of the distribution. Additional examples for young and intermediate-age stars are shown in Appendix C for comparison. The relation between the orange curve and the blue histogram shows how the recovered age posterior approximates the underlying distribution of ages in the training set. The histogram does not account for input uncertainties in the data, however the recovered posterior shown in orange does. This example happens to just barely succeed in recovering the age to $1\sigma$ and demonstrates how the posterior from normalizing flows can have asymmetric uncertainties.
@@
Figure 2: (Left) The larger black point marks the location of an example test-set star overlaid on the full training set (color-coded by [Fe/H]) shown in a Kiel diagram (Upper Left) and in the [C/N]–age relation (Lower Left). In the [C/N]–age plot, a red error bar indicates the model’s age estimate and uncertainty for the star. (Right) Age posterior for the example star. The black curve represents the full posterior computed with the input uncertainties, with the solid orange line indicating the maximum likelihood and dashed lines showing $1\sigma$ error bars. The red dashed line indicates the asteroseismic age, while the blue histogram displays the age distribution of training-set stars that are nearby in input parameter space. Additional example stars from the test set are shown in Appendix C
@@ 4.4 Model Predictions Beyond Stellar Ages
Figure 3: (Upper) 2D probability distribution of [C/Fe] & [N/Fe] for the star depicted in Fig 2. The red dot marks the maximum likelihood, and the red line indicates the $1\sigma$ boundary. The orange point and error bar shows the star’s measured [C/Fe] & [N/Fe] and associated uncertainties. The diagonals show lines of constant [C/N]. (Lower) Displays a similar plot for a different star with a higher [C/Fe] than the model predicts. Both plot exhibit slight banding features, these are most likely artifacts of the normalizing flow model over fitting to local noise or small-scale features.
@@ 4.5 Training Space Density
Figure 4: Parameter space coverage of stars with a training space density above $3\times 10^{9}$ (color-coded by density) compared to the full DR19 dataset (black points). These are scatter plots with each stars training density used for color coding. Regions with higher densities are better sampled in the training data, yielding more reliable age estimates. (Upper) Kiel diagram (Lower) [Fe/H] versus surface gravity ($\log g$). The sharp banded features seen in the black points are artifacts from the ASPCAP grid edges. This Figure is repeated in Appendix B.1 with different training space density thresholds.
We utilize this value as a measure of how well-sampled the parameter space around a given star’s parameters was during training, and we use it as a cutoff to limit age estimates to stars that occupy well-covered regions of parameter space. Figure 4 shows a Kiel diagram of the stars in our catalog, color-coded by their training space density. The right-hand panel shows the same sample but restricted to stars with a training space density greater than $3\times 10^{9}$, closely reflecting the parameter space coverage of the original training set. Choosing a smaller cutoff allows the model to extrapolate into less well-sampled regions, while a higher cutoff restricts the sample to stars within the most densely populated parts of parameter space. In subsequent sections, we refer to this value as the training space density, and we adopt a cutoff of greater than $3\times 10^{9}$ for the age validation. This choice demonstrates confidence in the model’s ability to extrapolate slightly beyond the asteroseismic core of the training data. We feel that this adopted value is a good starting point that is not too conservative, resulting in a small number of stars in the output, while also constraining the model from extrapolating too far. When applied to the training set, this value results in 95% of the training data being retained. Section 6.1 and Figure 11 further discuss the effects of this decision and alterations in star counts for the whole DR19 sample. Appendix B shows how varying the training space density used as a threshold changes the various plots and maps shown in this work.
@@ 5.1 Validation Against Asteroseismic ages
Figure 5 shows the maximum likelihood ages for the test set as determined by our model compared to the asteroseismic ages. The figure demonstrates that for ages $<8$ Gyr the model returns accurate ages. However, at ages $>8$ Gyr, the model’s maximum likelihood begins to diverge from the asteroseismic estimates. This result is not unexpected, as it has been observed and documented in multiple spectroscopic age studies (Leung & Bovy 2019; Mackereth et al. 2019; Anders et al. 2023; Stone-Martinez et al. 2024). The plateau at 8 Gyr is also expected from [C/N] ages, as shown in Roberts et al. (2024), due to mixing processes being weaker in lower mass (older) stars. Another potential contributor to the observed behavior is contamination of the training set by stars that have experienced mergers or mass transfer. Such stars can exhibit atypical [C/N] ratios for their true age. However, these objects constitute a small fraction of the training set and are unlikely to significantly shift the maximum likelihood estimates produced by the model.
@@ 5.2 Evaluating Model-Derived Uncertainties
Figure 7: (Upper) Histogram of the relative $1\sigma$ uncertainties for asteroseismic labels (blue), the averaged upper and lower relative $1\sigma$ uncertainties derived from the StarFlow model (orange), and the relative differences between the asteroseismic age and the StarFlow maximum likelihood ages (green). (Lower) Bar plot showing the median relative uncertainty recovered from the StarFlow age posteriors, binned by asteroseismic age.
@@ Acknowledgments
We are grateful to Juna Kollmeier, Jamie Tayar, and José G. Fernández-Trincado for valuable discussions and feedback. Much of the discussion in this manuscript was inspired by conversations with various conference and seminar attendees, including Phil Van-Lane, James Johnson, Jennifer Johnson, Marc Pinsonneault, Jason Jackiewicz, David Weinberg, Phillip Cargile, Henry Leung, Joshua Speagle, and Paul Beck.
A.S-M. gratefully acknowledges support from SDSS-V.
E.J.G. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2202135
Funding for the Sloan Digital Sky Survey V has been provided by the Alfred P. Sloan Foundation, the Heising-Simons Foundation, the National Science Foundation, and the Participating Institutions. SDSS acknowledges support and resources from the Center for High-Performance Computing at the University of Utah. SDSS telescopes are located at Apache Point Observatory, funded by the Astrophysical Research Consortium and operated by New Mexico State University, and at Las Campanas Observatory, operated by the Carnegie Institution for Science. The SDSS web site is www.sdss.org.
SDSS is managed by the Astrophysical Research Consortium for the Participating Institutions of the SDSS Collaboration, including the Carnegie Institution for Science, Chilean National Time Allocation Committee (CNTAC) ratified researchers, Caltech, the Gotham Participation Group, Harvard University, Heidelberg University, The Flatiron Institute, The Johns Hopkins University, L’Ecole polytechnique fédérale de Lausanne (EPFL), Leibniz-Institut für Astrophysik Potsdam (AIP), Max-Planck-Institut für Astronomie (MPIA Heidelberg), Max-Planck-Institut für Extraterrestrische Physik (MPE), Nanjing University, National Astronomical Observatories of China (NAOC), New Mexico State University, The Ohio State University, Pennsylvania State University, Smithsonian Astrophysical Observatory, Space Telescope Science Institute (STScI), the Stellar Astrophysics Participation Group, Universidad Nacional Autónoma de México, University of Arizona, University of Colorado Boulder, University of Illinois at Urbana-Champaign, University of Toronto, University of Utah, University of Virginia, Yale University, and Yunnan University.
This research was supported by the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP) which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy – EXC-2094 – 390783311.
@@ Appendix B Using a different training density threshold
Here we present how parameter space coverage (Figure 4), Galactic age map (Figure 10), and chemo-age map (Figure 13) change when varying the training density threshold used.
@@
B.1 Parameter space coverage of stars for different training density thresholds
@@
B.2 Galactic Age Map using different training density thresholds
@@
B.3 Chemo-Age map using different training density thresholds
@@
Appendix C Example posteriors
To complement Figure 2 in the main text, we include additional examples of test-set stars across a broader range of parameters and age shown in Figure 17.
[…]